Mr. Masterson has determined that the following items are not selling as much this time of year and he wants to order less of them. He wants to decrease the quantities he purchases by 2/3, ordering only 1/3 of the original amount. Please calculate the NEW amounts he needs to purchase based on the old amounts given.
The old amount is 160 lbs. of grapes so if I reduce that order by 2/3 what is the fractional answer?
but more than the answer I am also seeking the formula on how to solve for future questions

Found 2 solutions by Boreal, stanbon:
Answer by Boreal(15235)

It's easier to look at what he wants, which is 1/3 the original amount. That way, you don't have to subtract
It is 160/3, 160*(1/3)
If you subtract, it is 160-(160*2/3)=480/3-320/3=160/3.

Answer by stanbon(75887)

1): Mr. Masterson has determined that the following items are not selling as much this time of year and he wants to order less of them. He wants to decrease the quantities he purchases by 2/3, ordering only 1/3 of the original amount. Please calculate the NEW amounts he needs to purchase based on the old amounts given.
The old amount is 160 lbs. of grapes so if I reduce that order by 2/3 what is the fractional answer?
Ans: 160 - (2/3)160 = (1/3)160 = 53 1/3 ; round up to 54
-----------
Cheers,
Stan H.
----------